OpenStudy (briellshaw):

ILL MEDAL Evaluate the expression 3x2 - 8, when x = 4. A. 40 B. 48 C. 4 D. 15

pooja195 (pooja195):

3(4^2)=?

pooja195 (pooja195):

First start by figuring out what 4 x 4=? 4^2=4 x 4

OpenStudy (briellshaw):

16

pooja195 (pooja195):

good 3 x 16 - 8 =?

OpenStudy (briellshaw):

42

pooja195 (pooja195):

Try again... what's 16 x 3?

OpenStudy (briellshaw):

48

OpenStudy (briellshaw):

-8=40

pooja195 (pooja195):

3(4^2)−8 (3)(16)−8 =48−8

pooja195 (pooja195):

48-8=?

OpenStudy (briellshaw):

40

OpenStudy (anonymous):

\[3\ 4^2-8=40 \]

pooja195 (pooja195):

Well done =)

OpenStudy (briellshaw):

thx

pooja195 (pooja195):

yw

OpenStudy (anonymous):

good job